OpenStudy (anonymous):

a car traveling at a speed of 13 meters per second accelerates uniformly to a speed of 25 meters per second in 5.0 seconds. -a truck traveling at a constant speed covers the same total distance as the car in the same 5.0 second interval. Determine the speed of the truck. -Calculate the magnitude of the accelerate of the car during this 5.0 second interval.

OpenStudy (loser66):

The car: accelerate \[\large a = \dfrac{v_{final}-v_{initial}}{\triangle t} = \dfrac{25-13}{5}= 6 ~~m/s^2\] so, the distance it covers in 5 seconds is \[\large S = \dfrac{v_{final}^2-v_{initial}^2}{2a}=\dfrac{25^2-13^2}{12}=38~~m\] The truck: On the same distance with the car 38m, in 5seconds , its speed is \[\large v= \dfrac{distance}{time}=\dfrac{38}{5}=7.6~~m/s\]

OpenStudy (anonymous):

@Loser66 Your processes are correct, but methinks the arithmetic is a bit off ^_^ \[a = \frac{\Delta v}{\Delta t} = \frac{v-v_0}{t} \\ \ \\ \ \\ \ =\frac{25m/s - 13m/s}{5s}\] \[a = \frac{12}{5}m/s^2\] Then to find the distance traveled by the car, I'm gonna use the first equation of motion instead of the 4th (both are valid, of course) \[ s = v_0 t + \frac{1}{2}at^2 \\ =(13m/s)(5s) + \frac{1}{2}\left(\frac{12}{5}m/s^2\right)(5s)^2 \\ \ \\ = (13m/s)(5s)+\left(\frac{6}{5}m/s^2\right) ( 25s^2) \\ \ \\ \ \\ =(13m/s)(5s)+ (6m/s)(5s) \\ \ \\ s= 95m\] And then the velocity of the truck is given, as stated above by \[v = \frac{\Delta s}{\Delta t} = \frac{95m}{5s}\] \[ v = 19m/s\]

OpenStudy (loser66):

thanks for correcting me,